# FiT: Flexible Vision Transformer for Diffusion Model

### v1

@@ 4 Experiments
Method Train Cost 256$\times$256 (1:1) 160$\times$320 (1:2) 128$\times$384 (1:3) FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ BigGAN-deep - 6.95 7.36 171.4 0.87 0.28 - - - - - - - - - - StyleGAN-XL - 2.30 4.02 265.12 0.78 0.53 - - - - - - - - - - MaskGIT 1387k×$\times$256 6.18 - 182.1 0.80 0.51 - - - - - - - - - - CDM - 4.88 - 158.71 - - - - - - - - - - - - U-ViT-H/2-G (cfg=1.4) 500k×$\times$1024 2.35 5.68 265.02 0.82 0.57 6.93 12.64 175.08 0.67 0.63 196.84 95.90 7.54 0.06 0.27 ADM-G,U 1980k×$\times$256 3.94 6.14 215.84 0.83 0.53 10.26 12.28 126.99 0.67 0.59 56.52 43.21 32.19 0.30 0.50 LDM-4-G (cfg=1.5) 178k×$\times$1200 3.60 5.12 247.67 0.87 0.48 10.04 11.47 119.56 0.65 0.61 29.67 26.33 57.71 0.44 0.61 ${}^{{\dagger}}$ (cfg=3.8,s=4) 6500k×$\times$256 1.79 4.57 283.01 0.81 0.61 135.6 73.08 9.35 0.15 0.20 124.9 70.69 13.38 0.13 0.42 DiT-XL/2-G (cfg=1.50) 7000k×$\times$256 2.27 4.60 278.24 0.83 0.57 20.14 30.50 97.28 0.49 0.67 107.2 68.89 15.48 0.12 0.52 ${}^{\ast}$ (cfg=1.50) 1800k×$\times$256 4.27 9.99 249.72 0.84 0.51 5.74 10.05 190.14 0.74 0.55 16.81 20.62 110.93 0.57 0.52
Table 1: Benchmarking class-conditional image generation with in-distribution resolution on ImageNet dataset. “-G” denotes the results with classifier-free guidance. ${}^{{\dagger}}$: MDT-G adpots an improved classifier-free guidance strategy (Gao et al., 2023): $w_{t}=(1-\cos\pi(\frac{t}{t_{max}})^{s})w/2$, where $w=3.8$ is the maximum guidance scale and $s=4$ is the controlling factor. ${}^{\ast}$: FiT-XL/2-G adopts VisionNTK for resolution extrapolation.
Method Train Cost 320$\times$320 (1:1) 224$\times$448 (1:2) 160$\times$480 (1:3) FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ U-ViT-H/2-G (cfg=1.4) 500k×$\times$1024 7.65 16.30 208.01 0.72 0.54 67.10 42.92 45.54 0.30 0.49 95.56 44.45 24.01 0.19 0.47 ADM-G,U 1980k×$\times$256 9.39 9.01 161.95 0.74 0.50 11.34 14.50 146.00 0.71 0.49 23.92 25.55 80.73 0.57 0.51 LDM-4-G (cfg=1.5) 178k×$\times$1200 6.24 13.21 220.03 0.83 0.44 8.55 17.62 186.25 0.78 0.44 19.24 20.25 99.34 0.59 0.50 ${}^{{\dagger}}$ (cfg=3.8,s=4) 6500k×$\times$256 383.5 136.5 4.24 0.01 0.04 365.9 142.8 4.91 0.01 0.05 276.7 138.1 7.20 0.03 0.09 DiT-XL/2-G (cfg=1.50) 7000k×$\times$256 9.98 23.57 225.72 0.73 0.48 94.94 56.06 35.75 0.23 0.46 140.2 79.60 14.70 0.094 0.45 ${}^{\ast}$ (cfg=1.50) 1800k×$\times$256 5.42 15.41 252.65 0.81 0.47 7.90 19.63 215.29 0.75 0.47 15.72 22.57 132.76 0.62 0.47
Table 2: Benchmarking class-conditional image generation with out-of-distribution resolution on ImageNet dataset. “-G” denotes the results with classifier-free guidance. ${}^{{\dagger}}$: MDT-G adopts an aforementioned improved classifier-free guidance strategy. ${}^{\ast}$: FiT-XL/2-G adopts VisionNTK for resolution extrapolation. Our FiT model achieves state-of-the-art performance across all the resolutions and aspect ratios, demonstrating a strong extrapolation capability.
Arch. Pos. Embed. FFN Train 256$\times$256 (i.d.) 160$\times$320 (i.d.) 224$\times$448 (o.o.d.) FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ DiT-B Abs. PE MLP Fixed 44.83 8.49 32.05 0.48 0.63 91.32 66.66 14.02 0.21 0.45 109.1 110.71 14.00 0.18 0.31 Config A Abs. PE MLP Flexible 43.34 11.11 32.23 0.48 0.61 50.51 10.36 25.26 0.42 0.60 52.55 16.05 28.69 0.42 0.58 Config B Abs. PE SwiGLU Flexible 41.75 11.53 34.55 0.49 0.61 48.66 10.65 26.76 0.41 0.60 52.34 17.73 30.01 0.41 0.57 Config C Abs. PE + 2D RoPE MLP Flexible 39.11 10.79 36.35 0.51 0.61 46.71 10.32 27.65 0.44 0.61 46.60 15.84 33.99 0.46 0.58 Config D 2D RoPE MLP Flexible 37.29 10.62 38.34 0.53 0.61 45.06 9.82 28.87 0.43 0.62 46.16 23.72 35.28 0.46 0.55 FiT-B 2D RoPE SwiGLU Flexible 36.36 11.08 40.69 0.52 0.62 43.96 10.26 30.45 0.43 0.62 44.67 24.09 37.10 0.49 0.53
@@
Method 320$\times$320 (1:1) 224$\times$448 (1:2) 160$\times$480 (1:3) FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ FIDnormal-$\downarrow$ sFIDnormal-$\downarrow$ ISnormal-$\uparrow$ Prec.normal-$\uparrow$ Rec.normal-$\uparrow$ DiT-B 95.47 108.68 18.38 0.26 0.40 109.1 110.71 14.00 0.18 0.31 143.8 122.81 8.93 0.073 0.20 DiT-B + EI 81.48 62.25 20.97 0.25 0.47 133.2 72.53 11.11 0.11 0.29 160.4 93.91 7.30 0.054 0.16 DiT-B + PI 72.47 54.02 24.15 0.29 0.49 133.4 70.29 11.73 0.11 0.29 156.5 93.80 7.80 0.058 0.17 FiT-B 61.35 30.71 31.01 0.41 0.53 44.67 24.09 37.1 0.49 0.52 56.81 22.07 25.25 0.38 0.49 FiT-B + PI 65.76 65.45 29.32 0.32 0.45 175.42 114.39 8.45 0.14 0.06 224.83 123.45 5.89 0.02 0.06 FiT-B + YaRN 44.76 38.04 44.70 0.51 0.51 82.19 75.48 29.68 0.40 0.29 104.06 72.97 20.76 0.21 0.31 FiT-B + NTK 57.31 31.31 33.97 0.43 0.55 45.24 29.38 38.84 0.47 0.52 59.19 26.54 26.01 0.36 0.49 FiT-B + VisionYaRN 44.76 38.04 44.70 0.51 0.51 41.92 42.79 45.87 0.50 0.48 62.84 44.82 27.84 0.36 0.42 FiT-B + VisionNTK 57.31 31.31 33.97 0.43 0.55 43.84 26.25 39.22 0.48 0.52 56.76 24.18 26.40 0.37 0.49
@@ 4.4 FiT In-Distribution Resolution Results
Following our former analysis, we train our highest Gflops model, FiT-XL/2, for 1.8M steps. We conduct experiments to evaluate the performance of FiT at three different in distribution resolutions: $256\times 256$, $160\times 320$, and $128\times 384$. We show samples from the FiT in Fig FiT: Flexible Vision Transformer for Diffusion Model, and we compare against some state-of-the-art class-conditional generative models: BigGAN (Brock et al., 2018), StyleGAN-XL (Sauer et al., 2022), MaskGIT (Chang et al., 2022), CDM (Ho et al., 2022), U-ViT (Bao et al., 2023), ADM (Dhariwal & Nichol, 2021), LDM (Rombach et al., 2022), MDT (Gao et al., 2023), and DiT (Peebles & Xie, 2023). When generating images of $160\times 320$ and $128\times 384$ resolution, we adopt PI on the positional embedding of the DiT model, as stated in Section 4.3. EI is employed in the positional embedding of U-ViT and MDT models, as they use learnable positional embedding. ADM and LDM can directly synthesize images with resolutions different from the training resolution.

### v4

@@ 4 Experiments
[…]
Table 1: Benchmarking class-conditional image generation with in-distribution resolution on ImageNet dataset. “-G” denotes the results with classifier-free guidance. : MDT-G adpots an improved classifier-free guidance strategy (Gao et al., 2023): $w_{t}=(1-\cos\pi(\frac{t}{t_{max}})^{s})w/2$, where $w=3.8$ is the maximum guidance scale and $s=4$ is the controlling factor. : FiT-XL/2-G adopts VisionNTK for resolution extrapolation.
[…]
Table 2: Benchmarking class-conditional image generation with out-of-distribution resolution on ImageNet dataset. “-G” denotes the results with classifier-free guidance. : MDT-G adopts an aforementioned improved classifier-free guidance strategy. : FiT-XL/2-G adopts VisionNTK for resolution extrapolation. Our FiT model achieves state-of-the-art performance across all the resolutions and aspect ratios, demonstrating a strong extrapolation capability.
[…]
@@
[…]
@@ 4.4 FiT In-Distribution Resolution Results
Following our former analysis, we train our highest Gflops model, FiT-XL/2, for 1.8M steps. We conduct experiments to evaluate the performance of FiT at three different in distribution resolutions: $256\times 256$, $160\times 320$, and $128\times 384$. We show samples from the FiT in Fig 1, and we compare against some state-of-the-art class-conditional generative models: BigGAN (Brock et al., 2018), StyleGAN-XL (Sauer et al., 2022), MaskGIT (Chang et al., 2022), CDM (Ho et al., 2022), U-ViT (Bao et al., 2023), ADM (Dhariwal & Nichol, 2021), LDM (Rombach et al., 2022), MDT (Gao et al., 2023), and DiT (Peebles & Xie, 2023). When generating images of $160\times 320$ and $128\times 384$ resolution, we adopt PI on the positional embedding of the DiT model, as stated in Section 4.3. EI is employed in the positional embedding of U-ViT and MDT models, as they use learnable positional embedding. ADM and LDM can directly synthesize images with resolutions different from the training resolution.